# Perceptions of Linguistic Uncertainty by Language Models and Humans

### v1

@@ D.2 Summary Metrics
Table 4: Summary metrics averaged across uncertainty expressions. All metrics are computed with respect to the reference distribution in the non-verifiable setting (Human + NV). “PA” reports the general agreement between LLMs and the mode of the human distribution, reported in percentages. “MAE” reports the absolute error between the mean responses of LLMs and those of humans. Wasserstein-1 computes the distance between LLM and the baseline distribution.
@@
D.3 Proportional Agreement
Tables 5 and 6 report the proportional agreement (PA) metric discriminated by uncertainty expression in the non-verifiable and verifiable settings, respectively. The results are reported in the filtered pool of human participants.
Table 5: Proportional Agreement (PA) of models in the non-verifiable setting.
Humans Mode Humans Individual GPT-4o GPT-4 ChatGPT Llama3 Mixtral 8x22B Gemini DBRX Random
Avg PA 27.6 17.6 18.9 24.4 19.7 23.6 21.8 25.4 14.2 5.1
@@ D.3 Proportional Agreement
highly likely 34.6 22.5 17.4 22.1 16.1 8.3 23.9 30.0 7.8 6.5
very likely 28.7 17.7 19.1 19.5 13.6 14.4 15.1 21.9 14.8 5.1
probable 16.0 11.1 14.9 14.3 7.6 14.8 15.7 14.1 10.3 5.1
likely 20.2 12.9 16.5 16.5 8.2 19.7 16.5 18.1 8.7 5.0
somewhat likely 20.2 13.5 12.8 17.9 6.8 16.5 18.2 16.2 6.3 5.7
somewhat unlikely 22.3 14.0 18.0 19.3 21.8 22.3 18.4 19.9 15.6 4.9
uncertain 35.1 16.9 35.1 35.1 33.1 35.1 35.1 35.1 35.1 3.4
possible 18.1 10.7 6.5 15.4 14.2 15.0 8.6 15.5 15.6 5.1
unlikely 19.1 13.5 11.1 18.3 15.9 16.8 10.1 16.5 14.1 4.8
not likely 18.1 12.6 11.3 17.1 16.7 17.7 16.1 18.1 10.7 5.7
doubtful 19.7 11.7 14.2 12.9 13.2 17.5 16.4 19.7 11.8 5.8
very unlikely 38.8 23.4 27.8 37.9 26.5 38.8 22.3 36.9 11.5 3.0
highly unlikely 35.1 23.5 24.8 35.1 25.7 35.1 28.7 32.8 10.9 3.8
Table 6: Proportional Agreement (PA) of models in the verifiable setting.
Humans Mode Humans Individual GPT-4o GPT-4 GPT-3.5 Llama3 Mixtral 8x22B Gemini DBRX Random
Avg PA 27.6 16.7 15.2 22.1 15.3 18.8 18.6 21.3 13.6 5.1
@@
highly likely 34.6 21.5 14.7 22.0 13.0 17.4 24.1 22.8 15.5 6.5
very likely 28.7 17.7 17.2 17.2 10.1 14.9 14.9 17.6 12.7 5.1
probable 16.0 11.4 11.3 11.5 3.8 6.8 5.8 9.9 6.6 5.1
likely 20.2 12.7 12.7 12.4 7.3 10.7 9.2 11.3 6.4 5.0
somewhat likely 20.2 13.3 7.0 12.1 6.5 8.6 9.0 10.0 7.6 5.7
somewhat unlikely 22.3 13.0 15.0 18.8 15.4 17.3 11.8 15.4 14.7 4.9
uncertain 35.1 15.9 30.3 34.2 25.1 34.0 33.4 34.0 31.8 3.4
possible 18.1 9.4 6.1 10.2 5.8 4.0 4.6 6.7 6.9 5.1
unlikely 19.1 13.2 10.6 16.8 13.6 15.0 9.4 14.3 12.6 4.8
not likely 18.1 11.7 10.9 15.0 12.8 15.5 13.3 15.7 12.7 5.7
doubtful 19.7 10.0 11.6 11.5 6.2 15.2 13.8 16.4 11.9 5.8
very unlikely 38.8 21.8 18.8 34.1 24.2 35.9 22.3 36.2 12.8 3.0
highly unlikely 35.1 22.7 17.2 32.8 22.1 33.3 28.7 33.3 14.0 3.8
D.4 Mean Response
Figure 14 illustrates the mean rated probability metric discriminated by uncertainty expressions across the non-verifiable, as well as the true and false verifiable statements.
@@ D.4 Mean Response
(b) highly likely
(c) highly unlikely
(d) possible
(e) very likely
(f) very unlikely
(g) probable
(h) likely
(i) not likely
(j) unlikely
(k) somewhat likely
(l) somewhat unlikely
(m) uncertain
(n) doubtful
Figure 14: Mean response (and $95\%$ confidence intervals) of verifiable statements discriminated by truthfulness of the statement across all 14 evaluated uncertainty expressions.
Appendix E Generalization results
The creation of this dataset mirrors the procedure described in Section 4. We manually repurposed 200 question-answer pairs from AI2 Arc (100 from the easy set and another 100 from the challenge set). For every statement, the authors produce a true statement and a false statement using the available information about the correct and incorrect multiple choices. The final dataset consists of 200 true statements and 200 false statements.
Table 7: Difference between the empirical distributions of each model in the original and the generalization datasets. Differences are computed for each uncertainty expression in terms of the Wasserstein distance between the two distributions. Smaller values imply similar distributions. We observe minimal changes in models like GPT-4o, GPT-4, and Gemini and larger differences in models like OLMo (7B), Gemma (2B), and ChatGPT.
GPT-4o GPT-4 ChatGPT Gemini Llama3 (70B) Mixtral 8x7B Mixtral 8x22B OLMo (7B) Gemma (2B)
Avg 1.9 1.8 9.1 4.2 8.8 7.6 3.7 13.1 48.5
[…]
highly likely 1.5 1.3 12.8 5.1 17.1 11.2 2.2 18.9 55.1
very likely 1.6 1.8 10.6 7.0 16.0 13.3 1.8 18.4 53.6
likely 4.3 2.9 10.9 8.1 13.8 10.4 9.0 14.3 51.4
probable 3.2 3.2 17.2 5.8 14.4 16.0 10.6 12.8 50.7
somewhat likely 2.6 4.5 5.1 4.4 8.5 7.8 11.1 8.0 42.2
possible 4.8 3.4 14.6 7.5 11.6 11.2 6.6 15.0 48.6
uncertain 0.5 0.7 12.8 1.5 1.6 2.2 1.7 10.9 47.9
somewhat unlikely 2.3 0.4 3.6 2.0 2.1 4.2 1.6 14.0 34.2
unlikely 0.7 1.1 4.6 3.4 4.3 6.3 1.3 8.4 42.3
not likely 0.7 1.8 8.3 2.4 4.7 0.6 1.5 13.6 48.7
doubtful 0.2 1.5 14.6 2.5 5.1 6.6 1.9 24.7 47.1
very unlikely 1.2 0.8 2.1 1.4 2.6 1.1 0.2 3.1 50.7
highly unlikely 1.0 1.0 3.0 2.1 3.3 1.8 0.2 7.1 54.4
Figure 15: Mean numerical response (and $95\%$ confidence intervals) for verifiable statements across true and false statements averaged over the uncertainty expressions in the generalization set. Similarly to the main experiment, we observe large absolute differences between the mean numerical responses produced by LLMs in true statements and false statements. However, the magnitude of these changes is slightly lower than in the original experiment.
(a) almost certain
(b) highly likely
(c) highly unlikely
(d) possible
(e) very likely
(f) very unlikely
(g) probable
(h) likely
(i) not likely
(j) unlikely
(k) somewhat likely
(l) somewhat unlikely
(m) uncertain
(n) doubtful
Figure 16: Mean response (and $95\%$ confidence intervals) of verifiable statements across true and false statements for the 14 evaluated uncertainty expressions in the generalization set.
Appendix F Probabilistic decoding
Table 8 summarizes the different metrics in the non-verifiable setting, when using probabilistic decoding with OpenAI models, i.e., temperature = 1. The probability distributions are obtained using the methodology described in Appendix C.
Table 8: Summary metrics average across uncertainty expressions using probabilistic decoding (temperature=1) in both non-verifiable and verifiable settings.

### v2

@@ E.1 Mode- and Distribution-Matching Metrics
Table 11: Summary metrics averaged across uncertainty expressions for both NV and V settings. All metrics are computed with respect to the human distribution in the non-verifiable setting (Human+NV). “PA” reports the general agreement between LLMs and the mode of the human distribution, reported in percentages. “MAE” reports the absolute error between the mean responses of LLMs and those of humans. Wasserstein-1 computes the distance between LLMs and human distributions.
@@
E.1.1 Proportional Agreement
As described in Section 4.3, the proportional agreement (PA) metric gauges the overall agreement between an agent’s and a reference (population) distribution. We use the results of the human studies in the non-verifiable setting as our reference distribution (dubbed Human Mode) throughout the whole paper.
Tables 12 and 13 report the proportional agreement (PA) metric discriminated by uncertainty expression in the non-verifiable and verifiable settings, respectively. The values are reported with respect to the population-level human distribution described in Section 3, denoted Human+NV.
Table 12: Proportional Agreement (PA) score per uncertainty expression in the non-verifiable setting. The scores are with respect to the population-level human reference distribution (Human+NV).
OLMo (7B) Gemma (2B) LLama3 (70B) LLama3 (8B) ChatGPT GPT-4 GPT-4o Mixtral 8x22B Mixtral 8x7B
Methodology full full full full top-k top-k top-k sampling sampling
Average 12.2 8.6 25.1 18.8 20.5 25.0 19.1 23.9 22.3 22.0
Standard Deviation 13.6 8.8 12.8 14.2 12.5 13.1 8.8 13.5 13.0 12.5
@@ E.1.1 Proportional Agreement
doubtful 2.6 5.5 18.9 9.8 13.2 12.9 14.2 17.5 16.4 13.7
highly likely 20.5 1.0 18.0 14.9 16.1 22.1 17.4 8.3 23.9 21.5
highly unlikely 14.9 10.5 35.0 19.8 25.7 35.1 24.8 35.1 28.7 25.2
not likely 4.2 4.6 17.8 15.5 16.7 17.1 11.3 17.7 16.1 10.3
possible 6.7 14.3 15.3 7.7 14.2 15.4 6.5 15.0 8.6 14.8
probable 7.9 3.3 15.2 7.0 7.6 14.3 14.9 14.8 15.7 13.5
somewhat likely 9.5 2.7 16.5 4.7 6.8 17.9 12.8 16.5 18.2 15.4
somewhat unlikely 1.0 15.5 22.3 21.5 21.8 19.3 18.0 22.3 18.4 19.3
uncertain 6.2 34.0 35.1 35.1 33.1 35.1 35.1 35.1 35.1 35.1
unlikely 1.6 9.2 16.6 16.8 15.9 18.3 11.1 16.8 10.1 16.4
very likely 14.4 0.8 17.4 15.2 13.6 19.5 19.1 14.4 15.1 19.1
very unlikely 13.9 9.2 38.3 15.9 26.5 37.9 27.8 38.8 22.3 21.8
Table 13: Proportional Agreement (PA) score per uncertainty expression in the verifiable setting. The PA scores are with respect to the population-level human reference distribution (Human+NV).
OLMo (7B) Gemma (2B) LLama3 (8B) ChatGPT GPT-4 GPT-4o LLama3 (70B) Mixtral 8x22B Mixtral 8x7B Gemini
Methodology full full full top-k top-k top-k sampling sampling sampling sampling
Average 7.8 7.0 10.3 15.9 22.8 15.3 19.5 19.4 15.6 22.1
Standard Deviation 12.2 8.9 5.1 11.5 13.8 7.1 10.8 14.6 9.3 13.2
@@
doubtful 0.9 4.2 11.2 6.2 11.5 11.6 15.2 13.8 9.3 16.4
highly likely 18.5 0.9 11.1 13.0 22.0 14.7 17.4 24.1 12.8 22.8
highly unlikely 8.3 1.2 4.0 22.1 32.8 17.2 33.3 28.7 23.9 33.3
not likely 2.6 1.5 14.0 12.8 15.0 10.9 15.5 13.3 7.5 15.7
possible 1.2 14.8 3.5 5.8 10.2 6.1 4.0 4.6 6.0 6.7
probable 1.8 3.3 6.5 3.8 11.5 11.3 6.8 5.8 5.7 9.9
somewhat likely 1.7 3.9 7.2 6.5 12.1 7.0 8.6 9.0 9.7 10.0
somewhat unlikely 0.8 15.0 10.6 15.4 18.8 15.0 17.3 11.8 19.6 15.4
uncertain 0.0 32.4 14.2 25.1 34.2 30.3 34.0 33.4 31.0 34.0
unlikely 0.4 11.3 17.7 13.6 16.8 10.6 15.0 9.4 11.6 14.3
very likely 11.6 0.7 9.3 10.1 17.2 17.2 14.9 14.9 11.1 17.6
very unlikely 7.5 1.7 4.1 24.2 34.1 18.8 35.9 22.3 18.6 36.2
E.1.2 Mean Absolute Error
Proposed in Section 4.3, Mean Absolute Error (MAE) measures the average agreement across uncertainty expressions between an agent’s distribution and a reference (population) distribution. In this case, we also use the human results from the non-verifiable setting as our reference distribution throughout the paper.
E.1.3 Wasserstein Distance
We use one-dimensional Wasserstein distance as a measure of the similarity between two conditional distributions. We use scipy.stats.wasserstein_distance and provide the 22 bins for each conditional distribution (i.e., all 21 bins from $0,5,...100$ but also the -1) and the corresponding normalized counts that were used to estimate the conditional distributions. Two identical distributions will be assigned a Wasserstein-1 distance of 0, whereas two maximally distant distributions will be assigned a Wasserstein-1 distance of 101.
Table 11 summarizes the distributional differences between the LLM conditional distributions for both LLM+NV and LLM+V settings with respect to the reference distribution (estimated in Section 3 and denoted Human+NV). We observe that overall the obtained results correlate well with the results reported by the MAE metric.
Other uses of distributional differences in the verifiable setting: we compare how different the empirical distributions obtained in the main experiment differ with respect to the empirical distributions observed in the generalization experiment. These results are reported in Section F.
E.1.4 Differences in Mean Numerical Responses
Figures 13 and 16 show the mean numerical response for the two verifiable datasets, discriminated by model and uncertainty expression. Overall, both plots show evidence of the large perceptual differences exhibited by different models according to the truthfulness of the evaluated statements.
(a) “almost certain”
(b) ‘‘highly likely’’
(c) “highly unlikely”
(d) “possible”
(e) “very likely”
(f) “very unlikely”
(g) “probable”
(h) “likely”
(i) “not likely”
(j) “unlikely”
(k) “somewhat likely”
(l) “somewhat unlikely”
(m) “uncertain”
(n) “doubtful”
Figure 13: Mean numerical response for the verifiable statements discriminated by truthfulness of the statements. We observe that the differences in mean numerical responses differ by uncertainty expressions. Focusing on the opposing linguistic expressions (e.g., “unlikely” vs “likely”), the mean numerical response gap observed for models seems to be systematically larger for positive expressions (e.g., “likely”) than for negative expressions (e.g., “unlikely”).
E.2 Pairwise Correctness Analysis
In our paper, we find that LLMs exhibit a systematic “knowledge bias,” where the average numerical response is higher for true statements than for false statements. To provide statistical support to this claim, we run a one-sided paired Wilcoxon signed-rank test, comparing mean response between true and false statements. We perform a paired test, as a true and false statement were generated from each original question (e.g., “water’s chemical formula is H2O” and “carbon monoxide’s chemical formula is H2O” are paired). Overall, we find that the knowledge bias is statistically significant at $\alpha=0.05$ for all models except OLMo (7B) and Gemma (2B)—see Tables 14, 15, and 16.
Table 14: Analysis of the pairwise differences between true and false statements in the verifiable dataset. Using a one-sided Wilcoxon signed-rank test, we find that, for most models, numerical responses for true statements are statistically significantly larger than those for false statements (* = significant at $\alpha=0.05$).
Model Methodology Statistic p-value
ChatGPT top-k 31180.00 <0.0001*
GPT-4 top-k 34980.00 <0.0001*
GPT-4o top-k 36185.50 <0.0001*
Gemini sampling 25564.00 <0.0001*
LLama3 (70B) sampling 34979.00 <0.0001*
Mixtral 8x22B sampling 16051.00 <0.0001*
OLMo (7B) sampling 5309.50 <0.0001*
LLama3 (8B) full 17115.00 <0.0001*
OLMo (7B) full 5309.50 <0.0001*
Gemma (2B) full 786.50 0.9979
Table 15: Analysis of the pairwise differences between true and false statements in the AI2-ARC (Easy) dataset. Using a one-sided Wilcoxon signed-rank test, we find that, for most models, numerical responses for true statements are statistically significantly larger than those for false statements (* = significant at $\alpha=0.05$).
Model Methodology Statistic p-value
ChatGPT top-k 183290.50 <0.0001*
GPT-4 top-k 303749.50 <0.0001*
GPT-4o top-k 346533.00 <0.0001*
Gemini sampling 180349.50 <0.0001*
LLama3 (70B) sampling 250250.00 <0.0001*
Mixtral 8x22B sampling 129417.50 <0.0001*
OLMo (7B) sampling 111281.00 <0.0001*
Gemma (2B) sampling 46702.50 0.1148
Table 16: Analysis of the pairwise differences between true and false statements in the AI2-ARC (Challenge) dataset. Using a one-sided Wilcoxon signed-rank test, we find that, for most models, numerical responses for true statements are statistically significantly larger than those for false statements (* = significant at $\alpha=0.05$).
ChatGPT top-k 123271.00 <0.0001*
GPT-4 top-k 194138.00 <0.0001*
GPT-4o top-k 231543.50 <0.0001*
Gemini sampling 106837.00 <0.0001*
LLama3 (70B) sampling 135453.00 <0.0001*
Mixtral 8x22B sampling 99689.50 <0.0001*
OLMo (7B) sampling 48763.50 0.0333
Gemma (2B) sampling 34161.50 0.0191
E.3 Variability
In this section, we conduct a quantitative analysis of the variability of the LLMs’ empirical distributions. This analysis is motivated by the observation that, visually, the empirical distributions obtained via greedy decoding appear to be less diverse than the population-level human distributions. In addition to drawing comparisons between humans and LLMs, we also inspect how the observed variability changes with different decoding algorithms, namely, we consider a probabilistic decoding algorithm temperature=1.
Metric. As a measure of the variability of an empirical distribution, we consider the interquartile range (IQR) (Mohr et al., 2022), i.e., the difference between the 0.75 and 0.25 quantiles. [^15] Other metrics could also be used to study the variability of the distributions, such as the entropy. However, we leave further analysis for future work. Intuitively, a smaller IQR value implies that the models’ predictions tend to be concentrated in the same set of values, whereas a larger IQR value suggests a more uniform distribution. Because we have an empirical distribution for each uncertainty expression, we use the average IQR across uncertainty expressions to analyse the overall behavior.
Results. In the non-verifiable setting, we find that most models exhibit lower average IQR values than population-level human perceptions (see Table 17). In particular, the values reported for GPT-4, GPT-4o, and LLama3 (70B) are up to 13x smaller than the one reported for humans. These findings suggest that, when using greedy decoding, these LLMs are unable to replicate the diversity of human responses.
When considering the change in avg IQR values for probabilistic decoding in Table 17, we observe an increase in the spread for most models with respect to the greedy decoding values (+6.60 and +10.60 IQR increase on average for 2-shot and 0-shot, respectively). In fact, from the evaluated models, GPT-4 and LLama3 (70B) remain mostly insensitive to the change of decoding algorithm (between 0 and 1.20) whereas the IQR values for ChatGPT and GPT-4o increase +4.38 and +35.3, respectively. These results seem to suggest that for some models, the variability of the observed empirical distributions may be a function of the employed decoding algorithm.
Table 17: InterQuartile Range (IQR) of the empirical distributions averaged across all uncertainty expressions in the non-verifiable setting. Reported values include both histograms created using greedy decoding (temperature=0) and random decoding (temperature=1). While most models exhibit become more diverse when using random decoding, models like GPT-4 and LLama3 (70B) seem to be minimally affected by the change in decoding algorithm, suggesting that these models lead to less diverse results (when compared to humans).
2-shot prompt 0-shot prompt
Models Methodology Greedy Probabilistic Greedy Probabilistic
Humans 15.00
[…]
GPT-4 top-k 1.15 1.15 1.15 2.69
GPT-4o top-k 3.85 39.15 3.85 39.00
LLama3 (70B) full 1.92 3.08 0.77 1.54
LLama3 (8B) full 15.38 21.54 10.00 21.92
OLMo (7B) full 36.54 39.23 36.15
Gemma (2B) full 23.08 19.62 0.77 11.54
Mixtral 8x22B sampling 5.00 5.77
Mixtral 8x7B sampling 4.23 4.62
E.4 Histograms
(a) Humans
(b) Random
(c) Gemma (2B)
(d) OLMo (7B)
(e) Gemini
(f) GPT-4o
(g) ChatGPT
(h) GPT-4
(i) LLama3 (8B)
(j) LLama3 (70B)
(k) Mixtral 8x7B
(l) Mixtral 8x22B
Figure 14: Models (and Human) empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting (NV). The empirical distributions are estimated using the greedy decoding algorithm (temperature=1). Highlighted boxes represent the mode value for each expression. With the exception of OLMo (7B) and Gemma (2B), LLMs are able to replicate the overall population-level human behavior: assigning lower numerical responses to uncertainty expressions (e.g., “highly unlikely” and “unlikely”) and progressively higher numerical responses until reaching the certainty expressions (e.g., “almost certain”).
Appendix F Generalization results
The creation of this dataset mirrors the procedure described in Section 4. We manually repurposed 200 question-answer pairs from AI2-ARC (100 from the easy set and another 100 from the challenge set). For every statement, the authors produce a true statement and a false statement using the available information about the correct and incorrect multiple choices. The final dataset consists of 200 true statements and 200 false statements.
Table 18: Analysis of the distributional differences between the conditional distributions estimated using greedy (temperature=0) versus probabilistic decoding (temperature=1). We use Wasserstein-1 distance to report the distributional differences between each conditional distribution: maximally distant distributions exhibit a score of 101.
0-shot prompting 2-shot prompting
Model NV V NV V AI2-ARC (Easy) AI2-ARC (Challenge)
ChatGPT 3.28 3.40 3.32 2.75 2.87 3.08
GPT-4 0.31 0.52 0.22 0.84 0.49 0.78
GPT-4o 19.88 15.51 15.38 19.86 20.70 20.97
Figure 15: Mean numerical response discriminated by truthfulness of statements for 400 verifiable statements derived from the AI2-ARC dataset. The mean numerical responses produced by LLMs when evaluated in the context of true statements is significantly larger than when evaluated with the false statements. Although the numerical response gap is lower in magnitude than those observed in the verifiable experiment (see Figure 6), the observed gap is still statistically significant.
@@ Appendix F Generalization results
(b) “highly likely”
(c) “highly unlikely”
(d) “possible”
(e) “very likely”
(f) “very unlikely”
(g) “probable”
(h) “likely”
(i) “not likely”
(j) “unlikely”
(k) “somewhat likely”
(l) “somewhat unlikely”
(m) “uncertain”
(n) “doubtful”
Figure 16: Mean numerical response for verifiable statements discriminated by truthfulness of statements for the AI2-ARC dataset. While the observed gaps are smaller in magnitude than in the main experiment, there is still a significant difference between mean numerical responses depending on the truthfulness of the statements. We hypothesize that the observed magnitude differences between the two datasets may be explained by semantic differences between the two QA datasets used to curate the verifiable statements. Future work should investigate this behavior more closely.
Appendix G Probabilistic Decoding
The findings reported in the main paper mostly concern the conditional distributions that are estimated using greedy decoding algorithm. Despite being a common decision in analyses papers (Yona et al., 2024; Steyvers et al., 2024), the use of greedy decoding may not provide the full picture of model behavior (Ivgi et al., 2024; Holtzman et al., 2020). Thus, to ensure that our results are not a degenerate behavior only observed when using greedy decoding algorithms, we conduct an analysis considering a probabilistic decoding (temperature=1): instead of estimating the conditional distributions for each uncertainty expression using the arg-max predicted numerical response, we use the available probability information to estimate the conditional distributions.
Most of our analyses on the study of OpenAI models— ChatGPT, GPT-4, and GPT-4o—using the top-k approach. Using OpenAI models has the benefit of being more cost-efficient than the other approaches, i.e., it is less time-consuming and lower cost than the full probability methodology or the sampling-based approach. Moreover, even though theoretically the use of full probability information is better for estimating the conditional distributions, we found that the smaller evaluated open-source models (e.g., OLMo (7B) and Gemma (2B)) were extremely sensitive to the prompt formatting and, frequently resulted in negligible probability mass being assigned to integers in the range [0, 100].
(a) ChatGPT
(b) GPT-4
(c) GPT-4o
Figure 17: Absolute Difference between the models’ conditional probability distribution when estimated using greedy (temperature=0) versus probabilistic decoding (temperature=1). The results refer to the non-verifiable dataset. Gray rectangles indicate bins whose empirical distribution with the greedy information has non-zero value. Bins are colored blue if the bin’s estimated probability is larger when using the probabilistic approach and are red if the probability was larger using the greedy approach. While we observe that the decoding choice has a larger impact for ChatGPT and GPT-4o, we find that it has minimal impact in GPT-4’s distributions, suggesting that GPT-4’s predictions tend to be very confident in its predictions to begin with.
Table 19: Summary metrics averaged across uncertainty expressions for both NV and V settings when using probabilistic decoding (temperature=1). All metrics are computed with respect to the human distribution in the non-verifiable setting (Human+NV). “PA” reports the general agreement between LLMs and the mode of the human distribution, reported in percentages. “MAE” reports the absolute error between the mean responses of LLMs and those of humans. Wasserstein-1 computes the distance between LLMs and human distributions.